OpenStudy (anonymous):

please help!! 1 medal and fan Are the lines y=-x-4 and 5x+5y=20 perpendicular? Explain. No; their slopes are not equal Yes; the product of their slopes is -1. No; their slopes are equal. Yes; their slopes are equal.

OpenStudy (ivettef365):

first step is isolate the y on equation 5x +5y = 20, to find the slope

OpenStudy (ivettef365):

so you have 5y = -5x + 20, then divide all numbers by 5 and you get y = -x + 4

OpenStudy (ivettef365):

so what is the slope on both equations

OpenStudy (anonymous):

y=mx+b

OpenStudy (ivettef365):

correct, where m = slope

OpenStudy (anonymous):

-5

OpenStudy (ivettef365):

no, the equations are y = -x -4 and the other one we said it is y = -x + 4

OpenStudy (anonymous):

so -x

OpenStudy (anonymous):

so the answer is d

jimthompson5910 (jim_thompson5910):

y=-x-4 is the same as y = -1x-4 it is of the form y = mx+b where m = -1 and b = -4 m is the slope, so the slope of y = -1x-4 is -1

jimthompson5910 (jim_thompson5910):

and when two lines have the same slope, they are parallel lines (not perpendicular lines)

OpenStudy (anonymous):

ok thanks guy